# Sweep Angle Effects of Flow Over an Undulated Cylinder

### v1

@@ 3.1 Effect of Sweep Angle on Fluid Forces
The time averaged coefficient of drag and root mean square lift coefficient are reported in Figure 3 for the three geometries and five sweep angle values at $Re$ = 250. Focusing first on the two smooth geometries of the cylinder and ellipse, there is a clear decrease in forces with respect to increasing sweep angle. The cylinder’s initial drag coefficient of 1.31 decreases to 0.33 when swept at 60 degrees, whereas the streamlined ellipse has $C_{D}=0.89$ for perpendicular flow and $C_{D}=0.26$ at $\Lambda=60$ degrees. Both smooth geometries closely follow a descending cosine curve as theorized by the independence principle [37], plotted as dashed curves in Figure 3(a). The model works well for smooth geometries since the decrease in the perpendicular flow component is responsible for separation and regular vortex shedding, the prominent source of drag force.
At $\mathit{\Lambda}=0$, the whisker geometry exhibits strong drag reduction with respect to a cylinder ($37.9\%$) and a smaller reduction compared to an ellipse ($8.6\%$). However, as sweep angle increases, the magnitude of drag forces for the whisker geometry quickly merges with and becomes indistinct from the smooth ellipse at $\mathit{\Lambda}=30$ and beyond, with both aligning with the circular cylinder at $\mathit{\Lambda}=60$. Thus, drag coefficient due to the undulated surface is only reduced when ($\mathit{\Lambda}<30$).
@@
Figure 3: Time-averaged drag and root mean square lift coefficient values for the circular cylinder, ellipse, and undulated whisker geometry with respect to sweep angle at $\bm{Re=250}$. Included as dashed curves are predictions of the forces based on the perpendicular value scaled by cosine of velocity as theorized by the independence principle [37].
Observing unsteady lift amplitude in Figure 3(b), a circular cylinder and ellipse exhibit a large difference at $\mathit{\Lambda}=0$ with RMS coefficients of 0.27 and 0.11, respectively. As sweep is added, this difference narrows as their forces both decline. Contrasting with drag force, the prediction of lift forces is less accurately approximated by the independence principle. This is likely due to to a greater dependence on swept vortex shedding characteristics and spanwise effects flow effects that do not develop commensurately to the normal velocity component.
Unlike drag, the root mean square lift exhibits stark differences between the whisker and ellipse. For perpendicular flow, the undulated surface drastically reduces $C_{L,RMS}$ compared to a circular cylinder and ellipse alike, with a reduction of 93.5 and 84.2%. Unlike the behavior of the smooth geometries, the whisker $C_{L,RMS}$ increases with low sweep angles, rising from 0.02 at $\mathit{\Lambda}=0$ to 0.07 at $\mathit{\Lambda}=15$. This still maintains a significantly smaller amplitude than an ellipse with a reduction of 35.4%. The $C_{L,RMS}$ value remains below an ellipse until a sweep angle of $\mathit{\Lambda}=45$ at which point values for an ellipse and whisker are roughly equivalent. This suggests that a point of diminishing returns in the effectiveness of the modeled whisker-inspired undulations exists at or near this sweep angle. At $\mathit{\Lambda}=60$, the undulated surface $C_{L,RMS}$ is higher than the ellipse though still below the circular cylinder and ellipse model.
@@
Figure 4: Averaged force coefficient values for the cylinder, ellipse, and whisker geometries with respect to sweep angle at $\bm{Re=500}$ as well as theoretical values of cylinder and ellipse forces based on the independence principle.
Simulations are also performed at a higher Reynolds number of 500 where similar results in forces are observed in Figure 4. In agreement with $Re=250$, drag coefficient decreases with increasing sweep angle for all geometries. The circular cylinder and ellipse once again demonstrate disparate drag coefficients in perpendicular flow of $C_{D}=1.17$ and 0.80 and almost converged coefficients of 0.25 and 0.22 at $\mathit{\Lambda}=60$. Both sets of coefficients still follow the predicted decreasing curve of the independence principle, however the prediction of drag force is most accurate while lift force has a greater deviation particularly at the largest sweep angles.
At $Re=500$, the whisker geometry experiences an increase in the effective drag reduction from an ellipse. In perpendicular flow, in drag coefficient coefficient is reduced by 10.4% compared to a previous 8.6%. The drag suppression also persists to a greater extent at 7.4% for $\mathit{\Lambda}=15$ where, previously in Figure 3, drag forces were observed to more quickly merge with ellipse force values at this angle. At $\mathit{\Lambda}=30$, a similar effect as the $Re=250$ simulations returns as the whisker geometry reaches a similar drag coefficient as the ellipse up to $\mathit{\Lambda}=60$.
Lift forces at $Re=500$ on the smooth cylinder and ellipse both experience a general increase over their $Re=250$ counterpart. Values of $C_{L,RMS}=0.34$ for a cylinder and 0.24 for an ellipse at $\mathit{\Lambda}=0$ both decrease with added sweep. The most interesting observation is the stronger $C_{L,RMS}$ suppression of the undulated cylinder at $Re=500$. For the perpendicular flow orientation, the whisker’s lift RMS coefficient of 0.021 is a 91.2% decrease from an ellipse. Again, conversely to the smooth comparison geometries, the whisker geometry shows an increasing $C_{L,RMS}$ with the addition of sweep at low angles. Despite this, at $\mathit{\Lambda}=15$ the undulated surface is responsible for a 75.5% reduction in lift force from the ellipse, and 50.0% at $\mathit{\Lambda}=30$. Similarly to the $Re=250$ simulations, the ellipse and whisker converge at $\mathit{\Lambda}=45$ and again diverge at $\mathit{\Lambda}=60$.
@@
To understand the force observations of the undulated cylinder, Figure 5 describes the interaction between sweep angle and the relative cross section. While possessing the same mean chord length, thickness, and resulting mean aspect ratio as the smooth ellipse geometry, the piece-wise sinusoidal surfaces of the whisker geometry vary the effective value of these parameters with respect to spanwise location and orientation. Figure 5(c) displays the relative thickness, $T_{rel}$ and chord length $C_{rel}$ as a function of span, $\mathit{z}$. As shown in the $T_{rel}$ plot, thickness varies with location along the whisker surface, however its magnitude is unaffected by sweep angle. Conversely, relative chord length is dependent on orientation and location. As with a smooth geometry, the overall magnitude of chord length increases with sweep, though the shape of its periodic variation changes. With no sweep angle, the location of maximum chord length aligns with the minimum thickness and vice versa. This relationship gradually changes as sweep angle increases, where at approximately 30 degrees the chord length profile flattens and then reverses its original trend afterwards where maximum chord length and thickness now coincide. These trends are a result of the 180 phase difference between thickness and chord undulations as well as the smaller phase difference between leading and trailing undulations in chord length.
@@
Together, these two dimensions create hydrodynamically important effects on the relative aspect ratio of the undulated cylinder, and can explain trends in the forces. In Figure 5(d), the sweep angle-dependent aspect ratio of the whisker geometry is represented as a vertical range. It is widest at zero sweep angle as the undulation phase difference along the $z$-axis creates cross-sections of thicker, round cylinder-like geometry alternating with regions that have a longer streamlined cross-section. As sweep angle increases, the undulation phase difference relative to sweep is reduced and then reversed, creating an increasingly uniform cross-section across the span. The range of deviation from the mean reaches a minimum at approximately 45 degrees, converging with the ellipse geometry and notably approximating its flow response and resulting forces which converge between the two geometries at this point.
3.2 Time Averaged Velocity
Time-averaged velocity profiles near the whisker surface can provide insight into the three-dimensional features of the flow. Imposing a sweep angle introduces a spanwise freestream velocity component which alters flow characteristics of the undulated geometry in distinct ways from the smooth comparison geometries. Visible in Figure 6(a), a smooth ellipse demonstrates a separation point that is constant with respect to span and moves downstream moderately with increasing sweep angle.
@@ 3.2 Time Averaged Velocity
For the whisker geometry in Figure 6(b), streamlines of time-averaged velocity demonstrate the increasing spanwise flow with sweep angle as well as its effect on separation and recirculation near the trailing edge. For the perpendicular flow orientation, streamlines are redirected around areas of increased thickness and converge in regions of increased chord length. Trailing edge streamlines showing reverse flow also highlight a unique separation line that is localized to each wavelength of the whisker surface and creates distinct shedding regions divided in the vicinity of the streamwise undulation peak.
Upon the addition of swept flow, the spanwise flow initiated by geometry undulations is diluted but not eliminated. At 15 degrees, spanwise flow across the surface follows the freestream angle with minor redirection when traversing the streamwise and transverse undulations. Here, the separation line that delineated two distinct wake regions has disappeared and is replaced by a continuous separation region. Although now connected across the entire span, the influence of surface undulations is clear as separation occurs noticeably closer to the trailing edge in the regions of greatest chord length and minimal thickness relative to incoming flow. Recirculating flow near the trailing edge is now more biased in the swept freestream direction, however it still fully reverses in the streamwise direction in regions adjacent to the previously mentioned delayed separation.
This feature is gradually lost as sweep angle increases. For the swept 30, 45, and 60 degree configurations, the separation point moves slightly towards the trailing edge, however, its spanwise variation weakens and settles to a function of the spanwise relative chord length. Additionally, flow behind the separation line gradually transitions from fully reversed in the streamwise direction to roughly perpendicular and later becomes mostly spanwise, increasingly approximating the spanwise symmetric behavior of smooth geometry. As time-averaged flow becomes increasingly ellipse-like, the change in observed lift and drag coefficient values converges with the ellipse as well.
(a) Smooth Ellipse
(b) Undulated Cylinder
Figure 6: Surface streamlines of time averaged velocity $\overline{u}$.
@@ 3.3 Vortex Wake Effects
Having observed the impact of swept geometry on forces and basic flow measurements, another important feature of whisker geometry is its modification of shed vortices. The flow structures created by the surface undulations are markedly three-dimensional in comparison to the smooth geometry and this has a strong influence on the mixing and dissipation of energy in the wake. To visualize these wake structures, Figure 7 shows isosurfaces of $Q$-criterion for the perpendicular and sweep angle flow orientations of the smooth ellipse and whisker geometry at Reynolds number 250.
@@
Figure 7: Flow structures for $\bm{Re=250}$ visualized by isosurfaces of $\bm{Q}$-criterion colored by the sign of $Z$-vorticity.
As increasing sweep angle increases the spanwise velocity component, fluid rotation strength on the axis parallel to the span ($\omega_{z}$) diminishes. At $\mathit{\Lambda}=0$ where rotation is strongest relative to shearing motion in the wake, a larger $Q$ value captures wake structures effectively for the whisker and ellipse. However, as sweep increases, the same $Q$ value washes out the weaker wake structures. Thus the value of $Q$ is scaled relative to the perpendicular flow configuration.
@@
Looking first into the properties of the smooth ellipse, the salient features up to and including the $\mathit{\Lambda}=45$ orientation are the spanwise symmetry in shed vortices, decreasing shedding frequency with increasing sweep angle, and primarily two-dimensional behavior relative to the streamwise direction. At the largest sweep angle of $\mathit{\Lambda}=60$, vortices are shed at a shallower angle to incoming flow in a pattern noted for similar flows in prior literature [31].
@@
These observations contrast with the flow structures shown in the wake of the whisker geometry. The seal whisker wake at $\mathit{\Lambda}=0$ consists of two distinct shedding regions that correspond to adjacent wavelengths of the whisker surface. This breakup of vortices along the span is a result of spanwise momentum transport over the undulated surface and leads to a noticeable increase in streamwise vorticity as is visible in the wake [5]. Also of note is the alternating phase pattern with which adjacent wavelengths exhibit shedding. Prior literature demonstrated that the presence of this pattern may be beneficial to the reduction of fluid forces, but is just one effect in an ensemble of others that produce the resulting lower forces [34, 35]. With the addition of sweep angle, this alternating phase pattern disappears, though some streamwise vorticity with the same spanwise periodicity as the surface undulations is still observable in the wake. In the $\mathit{\Lambda}=15$ swept results, due to a directional bias from the added spanwise flow, the streamwise vortex structures are no longer equally sized and their appearance is sensitive to the choice of $Q$ threshold. Beyond this point at $\mathit{\Lambda}=30$ and $45$, shed vortices are similar to those created in the wake of the smooth ellipse. There exists some separation between adjacent wavelengths in the near wake that is merged into span-symmetric roller vortices downstream. The ellipse-like behavior of flow structures also mirrors the converging trends in forces for the middle range of sweep angle values between the two geometries. Finally, for $\mathit{\Lambda}=60$, the whisker geometry is observed to delay the ellipse behavior of peeling off vortex structures at a shallow angle and instead maintains the roller-type structures for a larger sweep. From the perspective of forces, this delayed transition in shedding behavior is less effective at lift suppression compared to the ellipse.

### v2

@@ 3.1 Effect of Sweep Angle on Fluid Forces
The time-averaged coefficient of drag and root mean square lift coefficient are reported in Figure 4 for the three geometries and five sweep angle values at $Re$ = 250. Focusing first on the two smooth geometries of the cylinder and ellipse, there is a clear decrease in forces with respect to increasing sweep angle. The cylinder’s initial drag coefficient of 1.31 decreases to 0.33 when swept at 60 degrees, whereas the streamlined ellipse has $C_{D}=0.89$ for perpendicular flow and $C_{D}=0.26$ at $\Lambda=60$ degrees. For smooth bluff bodies, the monotonic reduction in force with increasing sweep is a consequence of decreasing the span-perpendicular velocity component that dominates pressure drag and near-wake forcing. This sweep-induced attenuation provides a comparison against which the effects of whisker undulation-induced flow modification can be evaluated.
At $\mathit{\Lambda}=0$, the whisker geometry exhibits strong drag reduction with respect to a cylinder ($37.9\%$) and a smaller reduction compared to an ellipse (9.0%). However, as sweep angle increases, the magnitude of drag forces for the whisker geometry quickly merges with and becomes indistinct from the smooth ellipse at $\mathit{\Lambda}=30$ and beyond, with both aligning with the circular cylinder at $\mathit{\Lambda}=60$. Thus, drag coefficient reduction attributable to the undulated surface is only observed when ($\mathit{\Lambda}<30$).
@@
Figure 4: Time-averaged drag and root mean square lift coefficient values for the circular cylinder, ellipse, and undulated whisker geometry with respect to sweep angle at $\bm{Re=250}$. Solid curves are normalized by a constant freestream velocity while dashed lines represent normalizing by the span-perpendicular velocity component in each orientation.
Observing unsteady lift amplitude in Figure 4(b), a circular cylinder and ellipse exhibit a large difference at $\mathit{\Lambda}=0$ with RMS coefficients of 0.51 and 0.11, respectively. As sweep is added, this difference narrows as their forces both decline. This behavior reflects the same sweep-induced reduction observed in drag, where decreasing span-perpendicular flow weakens coherent vortex shedding and associated lift fluctuations for smooth geometries.
Unlike drag, the root mean square lift exhibits stark differences between the whisker and ellipse. For perpendicular flow, the undulated surface drastically reduces $C_{L,RMS}$ compared to a circular cylinder and ellipse alike, with a reduction of 97.0% and 86.1%. Unlike the behavior of the smooth geometries, the whisker $C_{L,RMS}$ increases with low sweep angles, rising from 0.02 at $\mathit{\Lambda}=0$ to 0.07 at $\mathit{\Lambda}=15$. This increase indicates a weakening of the suppression effects of whisker undulations with the introduction of sweep, as spanwise flow components disrupt the three-dimensional wake interactions responsible for lift reduction. This still maintains a significantly smaller amplitude than an ellipse with a reduction of 36.9%. The $C_{L,RMS}$ value remains below an ellipse until a sweep angle of $\mathit{\Lambda}=45$ at which point values for an ellipse and whisker are roughly equivalent. This convergence represents a crossover between two competing effects: the monotonic sweep-induced reduction in unsteady forcing and the diminishing influence of whisker-induced lift suppression with increasing sweep. At this point, there is a notable similarity in forces magnitudes arising from different wake structures between each geometry, and a likely marker of diminishing returns in the effectiveness of the modeled whisker-inspired undulations at or near this sweep angle. To underscore this, at $\mathit{\Lambda}=60$, the undulated surface $C_{L,RMS}$ is higher than the ellipse though still below the circular cylinder and ellipse model.
The dashed curves in Figure 4 represent the force coefficients normalized by the perpendicular velocity component. Cylinder drag forces collapse onto a nearly flat trend, suggesting that the perpendicular flow component dominates the sweep angle effects for this geometry. Here, $C_{D\perp}$ reflects the approximate validity of the classical independence principle for a cylinder where drag scales with the perpendicular velocity component. The whisker and ellipse, however, show a noticeable upward trend at sweep angles of $\mathit{\Lambda}=30-60$, indicating additional contributions to drag from geometry-specific wake interactions that are over-corrected. This demonstrates a departure from IP predictions, particularly for the whisker geometry’s initial increase in drag at 15 degrees which is not accounted for in models normalizing by perpendicular velocity. Lift coefficients for all geometries are more complex, indicating the effect of this renormalization is less significant than for drag forces, suggesting that lift is perhaps more sensitive to the three-dimensional wake dynamics explored in later sections.
@@
Figure 5: Averaged force coefficient values for the cylinder, ellipse, and whisker geometries with respect to sweep angle at $\bm{Re=500}$ normalized by a constant freestream velocity and alternatively the span-perpendicular velocity component in each orientation.
Simulations at Reynolds number 500 exhibit anticipated stronger fluctuations and less orderly wake structures than $Re=250$. Forces follow similar expected mean trends in sweep-induced responses, and these expectations help distinguish between changes due to Reynolds number and undulation-specific effects. Drag coefficients for all geometries again decrease with increasing sweep angle, however the whisker shows a slightly stronger effective drag suppression with a larger reduction from the ellipse in perpendicular flow (11.4% versus 9.0%) with suppression persisting at $\mathit{\Lambda}=15$. Perpendicular coefficients diverge further from expected collapse of truly two-dimensional, perpendicular velocity scaled flows at this Reynolds number, particularly for larger sweep angles that are similarly over-corrected. The most significant change occurs in lift forces. In addition to the decrease in $C_{L,RMS}$ for all geometries at this Reynolds number, the whisker exhibits an even stronger suppression in perpendicular flow at 90.8% relative to an ellipse. While whisker RMS lift increases with sweep angle, the magnitude of the suppression remains higher for longer as the whisker demonstrates lower lift forces at $\mathit{\Lambda}=15$ and $30$ before converging with an ellipse at $\mathit{\Lambda}=45$. This point represents a whisker lift maximum and reflects a similar observation of competing effects identified at $Re=250$. Here, degradation of whisker-induced suppression mechanisms reaches a balance with sweep-induced attenuation; however, at this higher Reynolds number, the former remained in greater effect over a broader range of sweep angles. The ellipse and whisker force magnitudes again demonstrate an interesting convergence despite differing wakes, then diverge at $\mathit{\Lambda}=60$ similarly to $Re=250$. Finally, sweep-dependent lift magnitudes are also poorly collapsed by velocity scaling in this regime as flow conditions increasingly violate assumptions of quasi-two-dimensional flow necessary for estimation using IP, as will be shown in later sections.
@@
To understand the force observations of the undulated cylinder, Figure 6 describes the interaction between sweep angle and the relative cross section. While possessing the same mean chord length, thickness, and resulting mean aspect ratio as the smooth ellipse geometry, the piece-wise sinusoidal surfaces of the whisker geometry vary the effective value of these parameters with respect to spanwise location and orientation. Figure 6(c) displays the relative thickness, $T_{rel}$ and chord length $C_{rel}$ as a function of spanwise location ($z$). As shown in the $T_{rel}$ plot, thickness varies with location along the whisker surface, however its magnitude is unaffected by sweep angle. Conversely, relative chord length is dependent on orientation and location. As with a smooth geometry, the overall magnitude of chord length increases with sweep, though the shape of its periodic variation changes. With no sweep angle, the location of maximum chord length aligns with the minimum thickness and vice versa. This relationship gradually changes as sweep angle increases, where at approximately 30 degrees the chord length profile flattens and then reverses its original trend afterwards where maximum chord length and thickness now coincide. These trends are a result of the 180 phase difference between thickness and chord undulations as well as the smaller phase difference between leading and trailing undulations in chord length.
@@
Together, these two dimensions create hydrodynamically important effects on the relative aspect ratio of the undulated cylinder, and can explain trends in the forces specific to the surface changes. In Figure 6(d), the sweep angle-dependent aspect ratio of the whisker geometry is represented as a vertical range. It is widest at zero sweep angle as the undulation phase difference along the $z$-axis creates cross-sections of thicker, round cylinder-like geometry alternating with regions that have a longer streamlined cross-section. As sweep angle increases, the undulation phase difference relative to sweep is reduced and then reversed, creating an increasingly uniform cross-section across the span. The range of deviation from the mean reaches a minimum at approximately 45 degrees, converging with the ellipse geometry and notably approximating its flow response and resulting forces which converge between the two geometries at this point.
3.2 Time-Averaged Streamlines
Time-averaged streamlines near the whisker surface provide insight into the three-dimensional features of the flow. The resulting streamlines represent the mean surface flow topology (rather than instantaneous vortex structures) and the identified separation features should be interpreted as mean-flow indicators rather than discrete shedding events. Imposing a sweep angle drives a redirection of near-wall flow in the spanwise direction rather than reversing purely in the streamwise direction, and separation does not always correspond to strong reversal of the body-fixed $x$ velocity component. This same effect, however, alters flow characteristics of the undulated geometry in distinct ways from the smooth comparison geometries.
@@ 3.2 Time-Averaged Streamlines
Observing a cropped segment of ellipse geometry at $Re=250$ in Figure 7(a), the surfaces demonstrate spanwise-uniform flow and a separation point that is constant across the span. Separation angles indicated in Table 1 for an ellipse show a small monotonic increase with sweep until $\mathit{\Lambda}=60$ where the flow transitions to a different shedding behavior. Shaded areas correspond to regions where wall shear stress is reversed in the body-fixed $x$ direction as a reverse flow indicator, and generally coincide with the separation points indicated by the mean streamlines. For the whisker geometry in Figure 7(b), streamlines demonstrate the increasing spanwise flow with sweep angle. For the perpendicular flow orientation, streamlines are redirected around areas of increased thickness and converge in regions of increased chord length. Trailing edge streamlines showing reverse flow also highlight a unique separation line that is localized to each wavelength of the whisker surface and creates distinct shedding regions divided in the vicinity of the streamwise undulation peak. These distinct separation cells are also visible in the disconnected regions of reverse wall shear stress, a unique phenomenon for this geometry noticeable later in the shedding phase of instantaneous flow structures.
With a sweep angle, the spanwise flow initiated by geometry undulations is diluted but not eliminated. At 15 degrees, the separation line that delineated two distinct wake regions has disappeared and is replaced by a continuous separation region, but the exact separation angle still varies across the span within a range noted in Table 1. This feature is gradually diminished as sweep angle increases. For the swept 30, 45, and 60 degree configurations, the separation angles increase while the range between minimum and maximum values decrease, as spanwise variation weakens and settles to a function of the spanwise relative chord length.
This progressive loss of strong, localized separation associated with undulations coincides with the observed convergence of lift fluctuations toward ellipse-like behavior. Also, at $\mathit{\Lambda}=60$, shaded reverse wall shear stress in the $x$-direction no longer entirely coincides with converging streamline-indicated separation, demonstrating a departure from typical shedding behavior due to three-dimensional spanwise flow at the extreme angles.
Time-averaged streamline flow topology trends are largely similar for an ellipse at $Re=500$ as demonstrated in Figure 7(c) with an anticipated separation angle reduction compared to $Re=250$. However, Table 1 indicates separation angle across sweep does not increase monotonically owing to increasingly three-dimensional instability effects. Unlike $Re=250$, the seal whisker at $\mathit{\Lambda}=0$ still experiences significant spanwise flow near the locations of maximum chord length. However, entirely distinct shedding cells are now absent which is consistent with the reduced spanwise segmentation of vortex structures shown in later instantaneous wake visualizations. The remaining whisker sweep angles behave in similar fashion to their prior $Re=250$ counterparts representing similar geometric effects despite the transition towards greater turbulence and instability.
(a) $Re=250$ Smooth Ellipse
(b) $Re=250$ Undulated Cylinder
(c) $Re=500$ Smooth Ellipse
(d) $Re=500$ Undulated Cylinder
Figure 7: Surface streamlines of time-averaged velocity $\overline{u}$ with shading representing regions of negative wall shear stress in the body-fixed $x$ direction.
Table 1: Flow separation angles, $\bm{\theta_{sep}}$, measured from the leading stagnation point to the zero crossing of the body-fixed $\bm{x}$-component of wall shear stress. For the whisker geometry, separation spans a band along the model where minimum and maximum values indicate the range which occurs at varying $\bm{z}$ locations along the span.
$Re=250$ $Re=500$
$\mathit{\Lambda}$ Ellipse () Whisker min. () Whisker max. () Ellipse () Whisker min. () Whisker max. ()
0 140.6 116.5 132.1 106.2 159.2
[…]
30 142.1 127.0 152.2 126.4 116.4 143.8
[…]
60 147.9 130.0 161.4 134.6 120.0 171.0
@@ 3.3 Vortex Wake Effects
The flow structures created by the surface undulations are markedly three-dimensional in comparison to the smooth geometry, strongly influencing mixing and dissipation of energy in the wake. Figure 8 shows isosurfaces of $Q$-criterion for the smooth ellipse and undulated cylinder at Reynolds number 250. The wake visualizations shown are representative snapshots selected at comparable points in the shedding cycle.
@@
Figure 8: Flow structures for $\bm{Re=250}$ visualized by isosurfaces of $\bm{Q}$-criterion colored by the sign of $\bm{z}$-vorticity. As sweep angle increases the fluid rotation strength on the axis parallel to the span ($\bm{\omega_{z}}$) diminishes, thus the value of $\bm{Q}$ visualized is lower as sweep increases.
@@
The salient features of the ellipse at $\mathit{\Lambda}\leq 45$ are the spanwise symmetry in shed vortices, decreasing shedding frequency, and two-dimensional behavior. At the largest sweep angle of $\mathit{\Lambda}=60$, vortices are shed at a shallower angle, almost perpendicular to the sweep angle, a pattern noted for similar flows in prior literature [lucor_effects_2003].
@@
In contrast, the seal whisker wake at $\mathit{\Lambda}=0$ consists of two distinct shedding regions that correspond to adjacent wavelengths of the whisker surface. This breakup of vortices along the span is a result of spanwise momentum transport over the undulated surface and leads to a noticeable increase in streamwise vorticity [lyons_wavelength_2022]. Also of note is the alternating phase pattern with which adjacent wavelengths exhibit shedding, a pattern that disappears with the addition of sweep. For $\mathit{\Lambda}=15$ small streamwise vortex structures still exist but are much smaller. At $\mathit{\Lambda}=30$ and $45$, shed vortices are similar to those of the smooth ellipse with residual wavelength-scale separation in the near wake merging into more span-symmetric roller vortices downstream. While visible artifacts of the whisker undulations remain, the primary mechanisms governing lift become increasingly dominated by flow components normal to the projected surface, leading to converging RMS lift coefficient values between the whisker and ellipse geometries. This convergence occurs despite qualitative differences in instantaneous wake structures, illustrating how similar global force statistics shown in Figure 4(a) can arise from distinct local flow dynamics. As sweep increases, the decreasing perpendicular flow component over an ellipse results in weaker roller vortex structures, whereas the whisker’s force-suppression capability, effective at disrupting spanwise coherence at low sweep, becomes increasingly outweighed by sweep-induced spanwise transport that favors re-merging of wake structures and ellipse-like force behavior. At $\mathit{\Lambda}=60$, the whisker geometry is observed to delay the ellipse behavior of peeling off vortex structures at a shallow angle and instead maintains the roller-type structures for a larger sweep. From the perspective of forces, this delayed transition in shedding behavior is less effective at lift suppression compared to the ellipse.
(a) $\mathit{\Lambda}=0,Q=0.6$
(b) $\mathit{\Lambda}=15,Q=0.6$
(c) $\mathit{\Lambda}=30,Q=0.4$
(d) $\mathit{\Lambda}=45,Q=0.32$
(e) $\mathit{\Lambda}=60,Q=0.16$
Figure 9: Isosurfaces of $\bm{Q}$-criterion colored by the sign of $\bm{z}$-vorticity for $\bm{Re=500}$. Due to increased vortex strength, higher $\bm{Q}$ values are displayed compared to those reported for $\bm{Re=250}$.
The wake structures for $Re=500$ are shown in Figure 9, demonstrating increased local turbulence with finer and more complex vortex structures. Additionally, spanwise instabilities characteristic of the flow regime are prevalent, particularly visible for the ellipse in the form of many streamwise finger-like structures along the span. However, many of the main features and trends observed at $Re=250$ are still present. At $\mathit{\Lambda}\leq 15$, the smooth geometry forms stronger and more compact vortices close to the trailing edge compared to the whisker, reflective of the higher RMS lift values. In contrast, the undulated cylinder sheds vortices further from the model surface. For ${\Lambda}=0$ there is a notable phase difference in the shedding of horseshoe-type vortex structures from adjacent wavelengths, similar to $Re=250$. Likewise, when ${\Lambda}>0$, there is increasing spanwise symmetry similar to the lower Reynolds number observations, causing the whisker geometry RMS lift to gradually approach that of the ellipse with the $\mathit{\Lambda}$ = 60 whisker geometry once again delaying the transition of the flow and maintaining the roller-like structures, reflecting in the larger RMS lift compared to the ellipse.
It is worth noting that other studies, e.g. [menon_spanwise_2022], have extended the use of the $Q$-criterion by incorporating both positive and negative $Q$ values in combination with force partitioning methods to enable a deeper dissection of the interplay between vortex dynamics and force generation. These methods highlight the potential for future work in the area of whisker-inspired geometry, particularly in relation to the forces observed as a result three-dimensional wake of this complex geometry.